OpenStudy (rayford):

solve 2sqrtx+7-10=-8

OpenStudy (lgbasallote):

\[2\sqrt{x+7} - 10 = -8\] add 10 to both sides \[\implies 2\sqrt{x+7} = -8 + 10\] combine like terms \[\implies 2\sqrt{x+7} = 2\] divide both sides by 2 \[\implies \sqrt{x+7} = 1\] square both sides \[\implies x + 7 = 1\] subtract 7 from both sides \[\implies x = 1 - 7\] combine like terms \[x = -6\] does that help?

OpenStudy (rayford):

wow...very good..

OpenStudy (rayford):

very easy to understand.

OpenStudy (lgbasallote):

wonderful ^_^

OpenStudy (rayford):

may you answer one more?

OpenStudy (rayford):

What is the equation of the line that passes through the point (0, 3) and is perpendicular to the line y = 2x + 3?

OpenStudy (rayford):

ive been stuck for hours..

OpenStudy (lgbasallote):

what is the slope of y = 2x + 3

OpenStudy (rayford):

3?

OpenStudy (lgbasallote):

y = mx + b m is the slope

OpenStudy (lgbasallote):

the slope is the coefficient of x

OpenStudy (rayford):

ok im with you.

OpenStudy (rayford):

i feel so stupid -__-

OpenStudy (lgbasallote):

so what's the slope?

OpenStudy (rayford):

2 :0?

OpenStudy (rayford):

the slope is 2

OpenStudy (lgbasallote):

right...so what is the slope of the perpendicular line?

OpenStudy (rayford):

3? may you solve this then let me review? that is my strongest way of learning

OpenStudy (lgbasallote):

the slope of the perpendicular line is negative reciprocal

OpenStudy (lgbasallote):

and im also kind of studying too lol so i cant do it in one go